A bike of mass 250 kg, travels at 120 km/h round a circular track of 110 m radius. Calculate the centripetal force required to maintain the car in its circular path.


1
Expert's answer
2021-06-01T10:47:53-0400

"F=mv^2/r=\frac{250\cdot120^2}{3.6^2\cdot110}=2525\ N"